[figure: 9.7 cm]

ANOMALOUS TRANSPORT AND THERMAL PROPERTIES OF THE SOLID SOLUTIONS U_{1−x}Th_{x}Cu_{5}M (M = Al, Ga, In and Sn)

R. Troć, V.H. Tran, D. Kaczorowski and A. Czopnik W. Trzebiatowski Institute of Low Temperature and Structure Research, Polish Academy of Sciences, P. O. Box 1410, 50-950 Wrocław, Poland

UCu_{5}M-type compounds are moderate heavy fermion systems with an antiferromagnetic order below T_{N} = 18 and 24 K for M = Al and In, respectively, and a ferrimagnetic order with T_{C} = 54 K for M = Sn. We present the results of magnetoresistance (MR) measurements as a function of temperature and applied magnetic field for the three compounds and for their solid solutions containing Th, i.e. U_{1−x}Th_{x}Cu_{5}M. Only in the case of UCu_{5}In and its Th substituted alloys a substantial positive contribution to the MR was observed at low temperatures, as for normal antiferromagnets. For the Al- and Sn-based alloys the MR at low temperatures is negative, and its field variation proves the existence of Kondo-type interactions. On Th-substitution the magnetic phase transition is shifted to lower temperature and simultaneuosly the MR becomes less negative. The field dependence of the MR of these alloys can be well described by the Coqblin-Schrieffer model. Interestingly, the heat capacity measurements have revealed a general tendency to increase the linear coefficient γ with magnetic dilution of a given UCu_{5}M compound by Th-substitution. This feature reflects an enlargement of the effective mass of conduction electrons with increase in the unit cell volume.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : PLENARY LECTURE

Corresponding author : R. Troć

Address for correspondence : W.Trzebiatowski Institute of Low Temperature and Structure Research, Polish Academy of Sciences, P.O.Box 1410, 50-950 Wrocław, Poland

Email address : troc@highscreen.int.pan.wroc.pl